• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Jack

    Log entry time 20:50:19 on July26,2006

    Entry number 175443

    keyword=Left Arm detectors

    The crystals have been removed. SC2 and S2 are back in place. The S0 motion seems to work. Gas started flowing to SC2 around 5:30PM. I am finished with the detectors. Hold any further problems till tomorrow morning.